Assume that a certain process has an 80% learning curve effect and the first unit took
2000hrs to produce.
Required:
a) Compute the number of hours required to produce the first 32 units
b) Compute the number of hours required to produce 32nd unit.
c) Assume that the wage rate is Kshs.100 per hour, compute the Labor cost of
producing the last 16 units
1
Expert's answer
2015-03-05T08:42:34-0500
Assume that a certain process has an 80% learning curve effect and the first unit took 2000 hrs to produce.
Required:
a) The number of hours required to produce the first 32 units is 2000*1^ln(0.8/2) + 2000*2^ln0.4 + ... + 2000*32^ln0.4 = 29,357.24 hours.
b) The number of hours required to produce 32nd unit is 2000*0.8^5 = 655.36 hours.
c) If the wage rate is Kshs.100 per hour, the Labor cost of producing the last 16 units is TC = 100*2000*16^ln(0.4) + 100*2000*17^ln0.4 + ... + 100*2000*32^ln0.4 = $1,233,617